OpenStudy (anonymous):

Suppose a section of the test has 50 multiple choice questions. Ron earns one point for each right answer and loses 1/4 point for each wrong answer. Ron answered 41 questions and his score was 34 3/4. How many questions did he answer correctly? how many questions did he answer incorrectly?

OpenStudy (anonymous):

Must use equations!

OpenStudy (anonymous):

Okay, so think of it this way. he answered 41 problems, and his score was only 34 3/4 he for each right answer he got a total of one point let's name this variable "c" and for each wrong answer he got 1/4 of a point taken away. Let's name this variable "w". Setting up an equation what could possibly be it from here?

OpenStudy (anonymous):

1c- 1/4w= 41 ???? i dont know i need help please

OpenStudy (anonymous):

please help @dread907

OpenStudy (anonymous):

Okay so your equation should be set up more like this: Points Possible = 50 Correct Answers = C Wrong Answers = W Points Earned equals = 34 1/4 50 - 34 1/4 equals 15 3/4 so he got 15 1/4 points taken away 15 divided by 4 equals what.

OpenStudy (anonymous):

3.75

OpenStudy (anonymous):

because there are 4 quarters in one. So since that equals 3.75 then + 1/4 equals 4 questions wrong

OpenStudy (anonymous):

but i needan equation set up

OpenStudy (anonymous):

Yes the equation was set up as we've done this. 50 - 34 1/4 = y(we never set a variable for this.) y divided by 4 = u(nor this) u + 1/4 = w

OpenStudy (anonymous):

i need with the variables c and w like i did

OpenStudy (anonymous):

like you did*

OpenStudy (anonymous):

@zepdrix can you help me?

OpenStudy (anonymous):

can you please help?

zepdrix (zepdrix):

sec reading it c:

OpenStudy (anonymous):

ok thanks!

zepdrix (zepdrix):

C + W = 41 <- the number of CORRECT answers and WRONG answers totals 41, because he answered 41 questions. 1C - (1/4)W = 34.75 <- The number of points is given by this equation, 1 point for a correct answer, -1/4 point for a wrong answer, and his points totaled 34 and 3/4. So we have a system of 2 equations, and 2 unknowns. shouldn't be too bad from here. I'm getting a weird number for C though, so I wanna make sure I have this set up correctly. hmm

zepdrix (zepdrix):

Wait what happens if he doesn't answer a question, does he lose any credit for that? Oh they don't care about that :D they just want C and W. hmm

OpenStudy (anonymous):

lol yeah it doesnt include that info

zepdrix (zepdrix):

Oh ok now the numbers are working out nicely :) I was just putting it into my calculator wrong a second ago. Yah these equations should work for us. From here, we can either apply, SUBSTITUTION, or, ELIMINATION. Which are you more comfortable with? :D

OpenStudy (anonymous):

substitution

zepdrix (zepdrix):

Messing with the first equation gives us W = 41 - C. Plugging this into the second equation gives us,\[\large C-\frac{1}{4}(W)=37.75 \qquad \rightarrow \qquad C-\frac{1}{4}(41-C)=37.75\]

zepdrix (zepdrix):

From here you should be able to find C without too much trouble ^^ Don't forget to DISTRIBUTE the -1/4 to each term inside the brackets. If you're still confused, you can let me know! c:

OpenStudy (anonymous):

ok thanks alot i will do it now. after this can you help me with 2 others? :( sorry for the trouble

OpenStudy (anonymous):

wait! isnt it 34.75?

OpenStudy (anonymous):

you put 37.75

zepdrix (zepdrix):

It is 34.75, yes! sorry about that c: typo

zepdrix (zepdrix):

Yah I'll try to answer them if I'm not too busy :D Throw an @zepdrix in if I'm being too lazy XD lol

OpenStudy (anonymous):

ok please please help me when i done with this one please please? i really need help :c

OpenStudy (anonymous):

c=36 w=5

OpenStudy (anonymous):

@zepdrix is that right??

zepdrix (zepdrix):

Hmmmm I came up with different numbers XD lemme check again.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\huge C-\frac{1}{4}(41-C)=34.75\]\[\huge C-\frac{41}{4}+\frac{1}{4}C=34.75\]\[\huge \frac{5}{4}C=45\] \[\huge C=36\] Yesss good job! I must've done it with the 37.75 the first time I tried it ^^

OpenStudy (anonymous):

lol ok thanks :) can i type the 2nd +3rd questions here please?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

mmmmmmm fine :3

OpenStudy (anonymous):

Sandi wants to use up 14 eggs and 12 cups of milk. How many quiches and bread pudding recipes can she make? RECIPE: Bread pudding: 5 cups milk and 3 eggs Quiche: 1 cup milk and 4 eggs

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hmmm this one I'm not so sure :( I'm having trouble setting up the equations. Hmmmm

OpenStudy (anonymous):

it gives a hint, (Write a system of equations that involves the number of quiches,x and the nimber of bread pudding y and the ingredients

zepdrix (zepdrix):

Hmm ok let's try something...

zepdrix (zepdrix):

Bah I dunno :c soz, i suck at word problems lol

OpenStudy (anonymous):

lol aw man :c ok ill just try my best what about this one? Ron has 30 nickels and dimes totaling 2.40. How many of each does he have?

zepdrix (zepdrix):

The total number of nickles and dimes adds up to 30 coins: \[\huge N+D=30\] The VALUE of those coins, 10 cents per dime, 5 cents per nickle, totals 2.40:\[\huge .05N+.10D=2.40\] So this our system.

zepdrix (zepdrix):

Similarly to the first problem, we should be able to use SUBSTITUTION quite effectively c:

OpenStudy (anonymous):

i set up those equations but i got weird answers :C

OpenStudy (anonymous):

i got n= 1.2 and y=28.8

zepdrix (zepdrix):

For the second equation, let's multiply both sides by 20 to get rid of those ugly decimals!\[\huge N+2D=48\]That might make it easier to work through :)

OpenStudy (anonymous):

wait can we use substitution ? :c

zepdrix (zepdrix):

Yes, Substitution is good, but it's just a good idea to get rid of the decimals before substituting c: It's not entirely necessary though. Substitute away! c:

OpenStudy (anonymous):

i did and i got those answers :C idk what i did wrong

zepdrix (zepdrix):

From the first equation we get,\[\huge N=30-D\] Plugging this into the second equation gives us,\[\huge (30-D)+2D=48\] \[\huge 30+D=48\] \[\huge D=48-30\]Understand what we did there? :D

OpenStudy (anonymous):

but wait how did you get those numbers? :o

OpenStudy (anonymous):

2d and 48?

zepdrix (zepdrix):

\[\huge .05N+.10D=2.40\]Multiplying both sides by 20 gives us,\[\huge 20(.05N+.10D)=(2.40)20\] \[\huge N+2D=48\]We did this to simplify the second equation, getting rid of the decimals so it's easier to work with.

zepdrix (zepdrix):

Confused? <:o

OpenStudy (anonymous):

i tthink i got it 12 nickels anmd 18 dimes?

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

yay thanks :D and are you sure you dont know how to solve the egg and milk problem?

OpenStudy (anonymous):

cause im so lost on that one :C